OpenStudy (anonymous):

Write An Equation Of Each Line In Standard Form With Integer Coefficient 1.Y=1/2X-2 2.Y=-7X-9 *DIFFERENT PROBLEMS *

OpenStudy (anonymous):

1. 2=1/2x-y (I think? I'm not positive.)

OpenStudy (anonymous):

2. 9=-7x-y

OpenStudy (anonymous):

Your Not Positive ?

OpenStudy (anonymous):

Someone else please check my work before confirming, I don't want to give a wrong answer ;-;

OpenStudy (anonymous):

lol its ok

jimthompson5910 (jim_thompson5910):

Standard form is `Ax+By=C` where A,B,C are whole numbers (most books make A to be a positive number)

jimthompson5910 (jim_thompson5910):

example of standard form 2x+3y = 10

OpenStudy (anonymous):

Ohh

OpenStudy (anonymous):

2 will be a and 3 will be b and 10 will be c

jimthompson5910 (jim_thompson5910):

for my example, yes

jimthompson5910 (jim_thompson5910):

the goal is to somehow transform y=1/2x-2 into the form Ax+By=C

OpenStudy (anonymous):

Ok Can You Help .

jimthompson5910 (jim_thompson5910):

first multiply every term by 2 to clear out that fraction

OpenStudy (anonymous):

1/2 * 2 =1

OpenStudy (anonymous):

2 x 2 =4

jimthompson5910 (jim_thompson5910):

doing that will turn y=1/2x-2 into 2y = x - 4

jimthompson5910 (jim_thompson5910):

now move the x over

OpenStudy (anonymous):

Subtract 2 From Both Sides

jimthompson5910 (jim_thompson5910):

you need to get the variable terms to one side

OpenStudy (anonymous):

So We Dont Subtract 2 from both sides

jimthompson5910 (jim_thompson5910):

more like subtract x from both sides

OpenStudy (anonymous):

2X?

OpenStudy (anonymous):

OH NVM I GET WHAT YOUR SAYING X FROM BOTH SIDES SO 1 FROM BOTH SIDES

jimthompson5910 (jim_thompson5910):

so we now have -x+2y = -4 that's technically in standard form where A = -1, B = 2, C = -4 but most books I've seen make A > 0. Multiply both sides by -1 to get x - 2y = 4

OpenStudy (anonymous):

OH

OpenStudy (anonymous):

jim ty b/c i don't know what im doing sorry :(

OpenStudy (anonymous):

So My End Result Should Of Been X-2Y=4

jimthompson5910 (jim_thompson5910):

yeah I'd go with x-2y = 4

OpenStudy (anonymous):

Ok What About the second problem ?

jimthompson5910 (jim_thompson5910):

same idea: get all of the variable terms to one side

jimthompson5910 (jim_thompson5910):

you don't have to worry about fractions this time

OpenStudy (anonymous):

lord i be confusing myself z

OpenStudy (anonymous):

-7 wll be a corect ?

jimthompson5910 (jim_thompson5910):

how do you move that -7x over? what must you do to both sides?

OpenStudy (anonymous):

subtract 7

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):